# THE SIEGEL MODULAR GROUP IS THE LATTICE OF

### v2

@@ 2.3.2 The statement of the volume formula
Theorem 2.2 ([23, Thm. 3.7]).
Let $\mathbb{G}$ be an absolutely simple, simply connected algebraic group of rank $n$ and type $C_{n}$ defined over a totally real number field $K$. Let $\mu_{\infty}$ denote the product measure on $\mathbb{G}_{\infty}$, where (for each infinite place $\upsilon$) the measure $\mu_{\upsilon}$ on $\mathbb{G}(K_{\upsilon})$ is described in [23, Sect. 1.3, Sect. 1.4]. Let $\Lambda$ be the principal arithmetic subgroup determined by a coherent collection $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$ of parahoric subgroups $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$. Then we have the formula
@@
Let $\upsilon$ be any finite place of $K$. Given that the exponents $m_{i}$ of $\mathbb{G}$ are $m_{i}=2i-1$ for $i=1,\ldots,n$ (see [24, Sect. 2.4]), we observe that
@@
where $q_{\upsilon}=\#\mathfrak{f}_{\upsilon}=N(\mathfrak{p}_{\upsilon})$ denotes the size of the residue field $\mathfrak{f}_{\upsilon}$ of $K$ at $\upsilon$ (or, equivalently, the norm of the prime ideal $\mathfrak{p}_{\upsilon}\subseteq\mathcal{O}_{K}$ associated with $\upsilon$). Indeed, this follows from [24, Sect. 2.4] and the fact that $\mathbb{G}$ is of type $C_{n}$ and thus without outer forms. Upon multiplying the right-hand side above over all finite places $\upsilon$ of $K$, one recognizes the reciprocal of the Euler product defining $\zeta_{K}(2)\cdots\zeta_{K}(2n)$.
As noted above, the quasisplit inner $K$-form $\mathcal{G}$ of $\mathbb{G}$ splits over $K$. Therefore $L=K$, and the factor involving $D_{L}/D_{K}^{\mathopen{}\mathclose{{}\left[L:K}\right]}$ vanishes. Moreover, since the number field $K$ is totally real, its number of (inequivalent) infinite places equals $\mathopen{}\mathclose{{}\left[K:{\bf Q}}\right]=d_{K}$, and hence the product over the infinite places of $K$ can be rewritten as claimed (cf. also [23, Rem. 3.8]). The claimed formula now follows from these observations and the fact that the Tamagawa number $\tau_{K}(\mathbb{G})$ is equal to $1$ (see [14]). ∎
@@ 3.1 Lower bounds in terms of $d_{K}$, $D_{K}$, and $n$
Let $\mathcal{T}$ denote the set of finite places $\upsilon$ of $K$ with the property that $P_{\upsilon}$ is not hyperspecial. (Equivalently, by (2.8), $\mathcal{T}$ is the set of all $\upsilon$ such that $\mathbb{G}$ does not split over $K_{\upsilon}$.) From the discussion in Section 2.3.1 we know that the types $\Theta_{\upsilon}$ of the parahorics $P_{\upsilon}$ are without symmetries. By [7, Lemma 6.3] we then have the bound
@@
Remark. In order to keep things as simple as possible, we will mainly use the weaker of the two estimates in Lemma 3.1 in the sequel, as it has the benefit that its right-hand side depends only on $d_{K}$ and $h_{K}$ and not on the unit groups. On the other hand, since the index $\mathopen{}\mathclose{{}\left[U_{K}^{+}:U_{K}^{2}}\right]$ can be computed rather easily for many particular number fields, we will make use of the full strength of the lemma once our other analyses have identified a small list of candidates for $K$.
@@
where the numerical bound for the infinite product follows from [13, Lemma 1]. ∎
@@ I. The cases $\bm{n\geqslant 4}$
Together, these claims imply that for any $n\geqslant 4$, the lattice $\Gamma$ must ”come from ${\bf Q}$” in the sense discussed earlier. Indeed, we know that the covolume of $\Gamma_{0}$ in $\mathrm{Sp}_{2n}({\bf R})$ is smaller than $1.83\cdot\Pi(n)$. If $n\geqslant 4$ and $K\neq{\bf Q}$, so that $\mathopen{}\mathclose{{}\left[K:{\bf Q}}\right]=d_{K}\geqslant 2$, the three claims of the lemma then imply that
@@
$n$ $\Pi(n)$
1 $2.5\times 10^{-2}$
2 $9.8\times 10^{-5}$
3 $1.9\times 10^{-7}$
4 $3.9\times 10^{-10}$
5 $1.5\times 10^{-12}$
6 $1.6\times 10^{-14}$
7 $6.6\times 10^{-16}$
8 $1.5\times 10^{-16}$
9 $2.2\times 10^{-16}$
10 $2.9\times 10^{-15}$
11 $4.1\times 10^{-13}$
12 $7.5\times 10^{-10}$
13 $2\times 10^{-5}$
14 $9.96$
$\vdots$ $\vdots$
By computing approximations of the first values of $\Pi(n)$ (see the table) and using Stirling’s formula, one can easily see that $\Pi(n)$ is increasing for $n\geqslant 14$. Since $f(n)$ is certainly also increasing for $n\geqslant 14$, we only need to choose $A$ such that (3.11) holds for $n=3,\ldots,14$. By using lower bounds for the values of $\Pi(n)$ and inspecting the inequalities (3.11) for these values of $n$, we find that $A$ satisfies the claim (b) provided that
@@ 13 $2\times 10^{-5}$
We now prove part (c). By using the approximation of $\Pi(4)$ given in the table, we see that this claim is satisfied if
@@ 4.1 Estimates of the parahoric factors
It is possible, using Bruhat–Tits theory, to describe the Levi components $\textbf{M}_{\upsilon}$ for each parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ explicitly, and hence to compute the exact values of $e^{\prime}(P_{\upsilon})$ with the help of (2.10). Conveniently, for the case of a $K$-form of $\mathrm{Sp}_{4}$, these computations already exist in the literature. Namely, we obtain from [8, Sect. 3, Table 2] that for any place $\upsilon<\infty$ of $K$, one either has $\textbf{M}_{\upsilon}=\mathrm{Sp}_{2}(\mathbb{F}_{q_{\upsilon}})\times\,{}^{2}O_{2}(\mathbb{F}_{q_{\upsilon}})=\mathrm{SL}_{2}(\mathbb{F}_{q_{\upsilon}})\times\,{}^{2}O_{2}(\mathbb{F}_{q_{\upsilon}})$ or $\textbf{M}_{\upsilon}=\mathrm{Sp}_{4}(\mathbb{F}_{q_{\upsilon}})$. (Here ${}^{2}O_{2}$ denotes the non-split, quasisplit orthogonal group.) By computing the orders of these groups and using (2.10) (cf. also [8, eq. (2.12)], one then has the two possibilities
$e^{\prime}(P_{\upsilon})=1\quad\text{or}\quad e^{\prime}(P_{\upsilon})=T(q_{\upsilon}):=\frac{q_{\upsilon}^{4}-1}{2(q_{\upsilon}+1)}.$
@@
We have now established that $e^{\prime}(P_{\upsilon})=1$ for all finite places $\upsilon$ of $K$, meaning that $\mathbb{G}$ splits at all finite places by (2.8). To conclusively rule out $K$, we note that since $42n=$ 4, our discussion in Section 2.2 shows that $\mathbb{G}$ must be a special unitary group over quaternion algebra $\mathbb{H}_{K}$. Moreover, if $\upsilon$ denotes any place of $K$, we recall from [7, Lemma 2.3] that $\mathrm{SU}(\mathbb{H}_{K_{\upsilon}};\tau;h)$ is ${K_{\upsilon}}$-split if and only if the quaternion algebra $\mathbb{H}_{K_{\upsilon}}$ splits; that is, if and only if there is an isomorphism defined over ${K_{\upsilon}}$ such that
@@
By [31, Thm. 14.6.1], we then conclude that, because $\mathbb{G}$ is split at all finite places (which means that $\mathbb{H}_{K}$ splits at all finite places), the number of infinite places where $\mathbb{H}_{K}$ is ramified (i.e non-split) is either 00 or $2$. It cannot be the case that $\mathbb{H}_{K}$ is ramified at both real places of $K$, since $\mathbb{G}$ splits at $\upsilon=\upsilon_{0}$. Hence $\mathbb{G}$ must even be split at the other real place $\upsilon_{1}$ of $K$. However, this is impossible, as we have already established, since then $\mathbb{G}(K_{\upsilon_{1}})=\mathrm{Sp}_{4}({\bf R})$ is not compact.
@@ 4.2 Identifying the lattice $\Gamma$
By the arguments up until this point, we know that for $n\geqslant 2$, a lattice $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ of minimal covolume comes from ${\bf Q}$; that is, $\mathbb{G}$ is defined over ${\bf Q}$, and $\Gamma$ is the normalizer in $\mathbb{G}_{\infty}$ of the principal arithmetic subgroup $\Lambda=\mathbb{G}({\bf Q})\cap\prod_{\upsilon<\infty}P_{\upsilon}$.
@@
We only need to justify the formula in the case of even rank. However, this follows immediately from [7, eq. (4.9)] once we split the product in the numerator into two products depending on the parity of the indexing variable $j$. ∎
@@
To complete the proof, all that remains is to show that $\Lambda=\Gamma=\Gamma_{0}$, where the last equality holds up to conjugation.
@@
such that $g_{\upsilon}P_{\upsilon}g_{\upsilon}^{-1}=\mathrm{Sp}_{2n}({\bf Z}_{\upsilon})$. Indeed, $\mathrm{GSp}_{2n}$ acts transitively on the hyperspecial parahoric subgroups of the symplectic group (cf. [30, Sect. 2.5]). Because of the topology on the restricted product $\mathbb{A}_{\bf Q}$ and the coherence of the collection $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$, we can assume that $g_{\upsilon}=I$ for all places $\upsilon$ except finitely many. Doing so, we thus obtain an element
@@
The class number of a split reductive group is at most the class number of any maximal split torus (cf. [22, Corollary to Thm. 8.11]). Since ${\bf Q}$ has class number $1$, it follows that $\mathrm{GSp}_{2n}$ has class number $1$ over ${\bf Q}$, and we have
@@
In other words, $h\Gamma h^{-1}=\mathrm{Sp}_{2n}({\bf R})\cap\Gamma_{0}=\Gamma_{0}$. Since $h\in\mathrm{GSp}_{2n}({\bf Q})$ and $\chi(h)>0$, the lattice $\Gamma$ is conjugate to $\Gamma_{0}$ by $h^{\prime}=\chi(h)^{-1/2}h\in\mathrm{Sp}_{2n}({\bf R})$. This concludes the proof of Theorem 1.1.

### v3

@@ 2.3.2 The statement of the volume formula
Theorem 2.2 ([25, Thm. 3.7]).
Let $\mathbb{G}$ be an absolutely simple, simply connected algebraic group of rank $n$ and type $C_{n}$ defined over a totally real number field $K$. Let $\mu_{\infty}$ denote the product measure on $\mathbb{G}_{\infty}$, where (for each infinite place $\upsilon$) the measure $\mu_{\upsilon}$ on $\mathbb{G}(K_{\upsilon})$ is described in [25, Sect. 1.3, Sect. 1.4]. Let $\Lambda$ be the principal arithmetic subgroup determined by a coherent collection $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$ of parahoric subgroups $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$. Then we have the formula
@@
Let $\upsilon$ be any finite place of $K$. Given that the exponents $m_{i}$ of $\mathbb{G}$ are $m_{i}=2i-1$ for $i=1,\ldots,n$ (see [26, Sect. 2.4]), we observe that
@@
where $q_{\upsilon}=\#\mathfrak{f}_{\upsilon}=N(\mathfrak{p}_{\upsilon})$ denotes the size of the residue field $\mathfrak{f}_{\upsilon}$ of $K$ at $\upsilon$ (or, equivalently, the norm of the prime ideal $\mathfrak{p}_{\upsilon}\subseteq\mathcal{O}_{K}$ associated with $\upsilon$). Indeed, this follows from [26, Sect. 2.4] and the fact that $\mathbb{G}$ is of type $C_{n}$ and thus without outer forms. Upon multiplying the right-hand side above over all finite places $\upsilon$ of $K$, one recognizes the reciprocal of the Euler product defining $\zeta_{K}(2)\cdots\zeta_{K}(2n)$.
As noted above, the quasisplit inner $K$-form $\mathcal{G}$ of $\mathbb{G}$ splits over $K$. Therefore $L=K$, and the factor involving $D_{L}/D_{K}^{\mathopen{}\mathclose{{}\left[L:K}\right]}$ vanishes. Moreover, since the number field $K$ is totally real, its number of (inequivalent) infinite places equals $\mathopen{}\mathclose{{}\left[K:{\bf Q}}\right]=d_{K}$, and hence the product over the infinite places of $K$ can be rewritten as claimed (cf. also [25, Rem. 3.8]). The claimed formula now follows from these observations and the fact that the Tamagawa number $\tau_{K}(\mathbb{G})$ is equal to $1$ (see [16]). ∎
@@ 3.1 Lower bounds in terms of $d_{K}$, $D_{K}$, and $n$
Let $\mathcal{T}$ denote the set of finite places $\upsilon$ of $K$ with the property that $P_{\upsilon}$ is not hyperspecial. (Equivalently, by (2.8), $\mathcal{T}$ is the set of all $\upsilon$ such that $\mathbb{G}$ does not split over $K_{\upsilon}$.) From the discussion in Section 2.3.1 we know that the types $\Theta_{\upsilon}$ of the parahorics $P_{\upsilon}$ are without symmetries. By [8, Lemma 6.3] we then have the bound
@@
Remarks.
1) We should point out that, although the result [8, Lemma 6.3] is stated and proved in the context of a different real form of the symplectic group, its proof makes no special use of the structure of ${\bf G}$ at the real places of the relevant number field $k$ (in the notation of [8]). Rather, the argument only relies on the structure at the non-archimedean places. As such, all types of parahorics in $\mathbb{G}(K_{\upsilon})$ can occur within the framework of [8, Lemma 6.3], and its analysis is therefore exhaustive in our situation as well.
2) In order to keep things as simple as possible, we will mainly use the weaker of the two estimates in Lemma 3.1 in the sequel, as it has the benefit that its right-hand side depends only on $d_{K}$ and $h_{K}$ and not on the unit groups. On the other hand, since the index $\mathopen{}\mathclose{{}\left[U_{K}^{+}:U_{K}^{2}}\right]$ can be computed rather easily for many particular number fields, we will make use of the full strength of the lemma once our other analyses have identified a small list of candidates for $K$.
3) For the convenience of readers unfamiliar with Galois cohomology, we wish to mention the paper [2] where a version of the bound (3.1) is derived in the special case $h_{K}=1$ using other techniques than those in [8] and [6]. In particular, it is demonstrated here how a change in the involved parahoric subgroups can change the index $\mathopen{}\mathclose{{}\left[\Gamma:\Lambda}\right]$.
@@
where the numerical bound for the infinite product follows from [15, Lemma 1]. ∎
@@ I. The cases $\bm{n\geqslant 4}$
Together, these claims imply that for any $n\geqslant 4$, the lattice $\Gamma$ must “come from ${\bf Q}$” in the sense discussed earlier. Indeed, we know that the covolume of $\Gamma_{0}$ in $\mathrm{Sp}_{2n}({\bf R})$ is smaller than $1.83\cdot\Pi(n)$. If $n\geqslant 4$ and $K\neq{\bf Q}$, so that $\mathopen{}\mathclose{{}\left[K:{\bf Q}}\right]=d_{K}\geqslant 2$, the three claims of the lemma then imply that
@@
By computing approximations of the first values of $\Pi(n)$ and using Stirling’s formula, one can easily see that $\Pi(n)$ is increasing for $n\geqslant 14$. Since $f(n)$ is certainly also increasing for $n\geqslant 14$, we only need to choose $A$ such that (3.11) holds for $n=3,\ldots,14$. By using lower bounds for the values of $\Pi(n)$ and inspecting the inequalities (3.11) for these values of $n$, we find that $A$ satisfies the claim (b) provided that
@@
We now prove part (c). By using the approximation $\Pi(4)\approx 3.9\cdot 10^{-10}$, we see that this claim is satisfied if
@@ 4.1 Estimates of the parahoric factors
It is possible, using Bruhat–Tits theory, to describe the Levi components $\textbf{M}_{\upsilon}$ for each parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ explicitly, and hence to compute the exact values of $e^{\prime}(P_{\upsilon})$ with the help of (2.10). Conveniently, for the case of a $K$-form of $\mathrm{Sp}_{4}$, these computations already exist in the literature. Namely, we obtain from [9, Sect. 3, Table 2] that for any place $\upsilon<\infty$ of $K$, one either has $\textbf{M}_{\upsilon}=\mathrm{Sp}_{2}(\mathbb{F}_{q_{\upsilon}})\times\,{}^{2}O_{2}(\mathbb{F}_{q_{\upsilon}})=\mathrm{SL}_{2}(\mathbb{F}_{q_{\upsilon}})\times\,{}^{2}O_{2}(\mathbb{F}_{q_{\upsilon}})$ or $\textbf{M}_{\upsilon}=\mathrm{Sp}_{4}(\mathbb{F}_{q_{\upsilon}})$. (Here ${}^{2}O_{2}$ denotes the non-split, quasisplit orthogonal group.) By computing the orders of these groups and using (2.10) (cf. also [9, eq. (2.12)], one then has the two possibilities
$e^{\prime}(P_{\upsilon})=1\quad\text{or}\quad e^{\prime}(P_{\upsilon})=T(q_{\upsilon}):=\frac{q_{\upsilon}^{4}-1}{2(q_{\upsilon}+1)}.$ (4.3)
@@
We have now established that $e^{\prime}(P_{\upsilon})=1$ for all finite places $\upsilon$ of $K$, meaning that $\mathbb{G}$ splits at all finite places by (2.8). To conclusively rule out $K$, we note that since $42n=$ 4, our discussion in Section 2.2 shows that $\mathbb{G}$ must be a special unitary group over quaternion algebra $\mathbb{H}_{K}$. Moreover, if $\upsilon$ denotes any place of $K$, we recall from [8, Lemma 2.3] that $\mathrm{SU}(\mathbb{H}_{K_{\upsilon}};\tau;h)$ is ${K_{\upsilon}}$-split if and only if the quaternion algebra $\mathbb{H}_{K_{\upsilon}}$ splits; that is, if and only if there is an isomorphism defined over ${K_{\upsilon}}$ such that
@@
By [33, Thm. 14.6.1], we then conclude that, because $\mathbb{G}$ is split at all finite places (which means that $\mathbb{H}_{K}$ splits at all finite places), the number of infinite places where $\mathbb{H}_{K}$ is ramified (i.e non-split) is either 00 or $2$. It cannot be the case that $\mathbb{H}_{K}$ is ramified at both real places of $K$, since $\mathbb{G}$ splits at $\upsilon=\upsilon_{0}$. Hence $\mathbb{G}$ must even be split at the other real place $\upsilon_{1}$ of $K$. However, this is impossible, as we have already established, since then $\mathbb{G}(K_{\upsilon_{1}})=\mathrm{Sp}_{4}({\bf R})$ is not compact.
@@ 4.2 Identifying the lattice $\Gamma$
By the arguments up until this point, we know that for $n\geqslant 2$, a lattice $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ of minimal covolume comes from ${\bf Q}$; that is, $\mathbb{G}$ is defined over ${\bf Q}$, and $\Gamma$ is the normalizer in $\mathbb{G}({\bf R})$ of the principal arithmetic subgroup $\Lambda=\mathbb{G}({\bf Q})\cap\prod_{\upsilon<\infty}P_{\upsilon}$.
@@
We only need to justify the formula in the case of even rank. However, this follows immediately from [8, eq. (4.9)] once we split the product in the numerator into two products depending on the parity of the indexing variable $j$. ∎
@@
To complete the proof of Theorem 1.1, all that remains is to show that $\Lambda=\Gamma=\Gamma_{0}$, where the last equality holds up to conjugation.
@@
such that $g_{\upsilon}P_{\upsilon}g_{\upsilon}^{-1}=\mathrm{Sp}_{2n}({\bf Z}_{\upsilon})$. Indeed, $\mathrm{GSp}_{2n}$ acts transitively on the hyperspecial parahoric subgroups of the symplectic group (cf. [32, Sect. 2.5]). Because of the topology on the restricted product $\mathbb{A}_{\bf Q}$ and the coherence of the collection $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$, we can assume that $g_{\upsilon}=I$ for all places $\upsilon$ except finitely many. Doing so, we thus obtain an element
@@
The class number of a split reductive group is at most the class number of any maximal split torus (cf. [24, Corollary to Thm. 8.11]). Since ${\bf Q}$ has class number $1$, it follows that $\mathrm{GSp}_{2n}$ has class number $1$ over ${\bf Q}$, and we have
@@
In other words, $h\Gamma h^{-1}=\mathrm{Sp}_{2n}({\bf R})\cap\Gamma_{0}=\Gamma_{0}$ where $h\in\mathrm{GSp}_{2n}({\bf Q})$. Taking into account the isomorphism (now, automorphism) $\mathrm{Sp}_{2n}({\bf R})\xrightarrow{\sim}\mathbb{G}({\bf R})=\mathrm{Sp}_{2n}({\bf R})$ induced by the distinguished real place $\upsilon_{0}$ of $K={\bf Q}$, we conclude that $\Gamma$ is conjugate to $\Gamma_{0}$ by an element of $\mathrm{GSp}_{2n}({\bf R})$ (cf. [11]). This concludes the proof of Theorem 1.1.
4.3 A final remark on non-special parahorics
We have proved Theorem 1.1 under the assumption that every parahoric had maximal volume. We will now justify this assumption, as promised in Section 2.3.1.
For the remainder of the paper, we will deviate from our previous notation and use the symbols $\Gamma$ and $\Lambda$ as free variables that denote, respectively, any maximal lattice in $\mathrm{Sp}(2n,{\bf R})$ and the principal arithmetic subgroup it normalizes (by the maximality criterion [28, Satz 3.5] of Rohlfs). Likewise, $\mathbb{G}$ will denote the algebraic $K$-group that defines $\Lambda$. $P_{\upsilon}(\Lambda)=P_{\upsilon}(\Gamma)$ will denote the parahoric subgroup at $\upsilon$ associated to the pair $(\Lambda,\Gamma)$. Finally, we will also write $\Lambda^{\mathrm{max}}$ to denote a principal arithmetic subgroup for which every $P_{\upsilon}(\Lambda)$ is of maximal volume; and $\Gamma^{\mathrm{max}}$ will then denote the normalizer of $\Lambda^{\mathrm{max}}$. In the same vein, for a pair $(\Lambda,\Gamma)$, we will write $\Lambda^{\mathrm{max}}$ and $\Gamma^{\mathrm{max}}$ to denote the corresponding “maximized” principal arithmetic subgroup (as explained in [5, Sect. 4.3] and [6, Sect. 3.8]) and its normalizer in $\mathbb{G}({\bf R})$, respectively.
Up until this point, we have proved the following result in particular.
Theorem 4.2.
Let $V_{0}$ denote the minimal covolume of all lattices in $\mathrm{Sp}(2n,{\bf R})$. If a lattice of the form $\Gamma^{\mathrm{max}}$ has covolume $V_{0}$, then $\mathbb{G}$ splits at every finite place of $K$, and every parahoric $P_{\upsilon}$ associated to $\Gamma^{\mathrm{max}}$ satisfies $e^{\prime}(P_{\upsilon})=1$.
We now prove that any lattice of minimal covolume must, in fact, be of the form $\Gamma^{\mathrm{max}}$.
Lemma 4.3.
Suppose that $\Gamma\subseteq\mathrm{Sp}(2n,{\bf R})$ is a maximal lattice, associated to an algebraic $K$-group $\mathbb{G}$, with covolume $V_{0}$. Then $\Gamma=\Gamma^{\mathrm{max}}$; that is, every parahoric $P_{\upsilon}(\Gamma)$ has maximal volume among all parahoric subgroups of $\mathbb{G}(K_{\upsilon})$.
Proof.
The discussion in Section 2.3.1 shows that also $\Gamma^{\mathrm{max}}$ has covolume $V_{0}$. Since both $\Gamma$ and $\Gamma^{\mathrm{max}}$ come from $\mathbb{G}$, Theorem 4.2 implies that $\mathbb{G}$ splits at all finite places — in particular, a parahoric subgroup of $\mathbb{G}(K_{\upsilon})$ is special if and only if it is hyperspecial, and consequently $\Gamma=\Gamma^{\mathrm{max}}$ if and only if every $P_{\upsilon}(\Gamma)$ is special. Theorem 4.2 also gives $e^{\prime}(P_{\upsilon}(\Gamma^{\mathrm{max}}))=1$, so we obtain from Theorem 2.2 and the inequality [5, 4.3.(15)] that
$1=\frac{\mu(\mathbb{G}({\bf R})/\Gamma)}{\mu(\mathbb{G}({\bf R})/\Gamma^{\mathrm{max}})}=\frac{\mathopen{}\mathclose{{}\left[\Gamma^{\mathrm{max}}:\Lambda^{\mathrm{max}}}\right]}{\mathopen{}\mathclose{{}\left[\Gamma:\Lambda}\right]}\prod_{\upsilon<\infty}e^{\prime}(P_{\upsilon}(\Gamma))\geqslant\prod_{\upsilon<\infty}\frac{e^{\prime}(P_{\upsilon}(\Gamma))}{\#\Xi_{\Theta_{\upsilon}}},$ (4.4)
where $\Xi_{\Theta_{\upsilon}}\subseteq\mathrm{Aut}(\Delta_{\upsilon})$ is the subgroup of all diagram automorphisms that preserve the type $\Theta_{\upsilon}=\Theta_{\upsilon}(P_{\upsilon}(\Gamma))$ of $P_{\upsilon}(\Gamma)$, and $\Delta_{\upsilon}$ denotes the local Dynkin diagram of $\mathbb{G}$ over $K_{\upsilon}$.
Concretely, we have #$\#\Xi_{\Theta_{\upsilon}}\in\mathopen{}\mathclose{{}\left\{1,2}\right\}$ for every $\upsilon$, and in fact #$\#\Xi_{\Theta_{\upsilon}}=1$ when $P_{\upsilon}(\Gamma)$ is special (see [8, Sect. 3.1] or [6, Sect. 3.2]). Therefore, any factor corresponding to a special parahoric on the right-hand side of (4.4) must be equal to $1$. The crucial fact, as we will see below, is that the converse holds as well: If $P_{\upsilon}(\Gamma)$ is not special, then the corresponding factor exceeds $1$. As this is impossible, the proof will be concluded once we demonstrate this claim. Equivalently, we must show that $e^{\prime}(P_{\upsilon}(\Gamma))>2$ whenever $P_{\upsilon}(\Gamma)$ is not special.
Let $\upsilon$ be any fixed place such that $P_{\upsilon}(\Gamma)$ is not special. If the residue field of $K_{\upsilon}$ has degree $q_{\upsilon}=2$ and $\mathbb{G}$ has rank $n=2$, then our claim follows immediately from the inequality
$e^{\prime}(P_{\upsilon}(\Gamma))\geqslant\frac{q_{\upsilon}^{4}-1}{2(q_{\upsilon}+1)}=\frac{15}{6}>2,$
which follows by inspection of [9, Sect. 3, Table 2]. We therefore have either $n\geqslant 3$ or $q_{\upsilon}\geqslant 3$. In this situation, we can use the “volume rigidity estimate” proved for non-special parahorics in [25, Prop. 2.10.(iv)] together with (2.13) (which is independent of the assumption we are in the process of justifying) to conclude that
$e^{\prime}(P_{\upsilon}(\Gamma))\geqslant\frac{q_{\upsilon}^{n+1}}{q_{\upsilon}+1}\prod_{j=1}^{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{q_{\upsilon}^{2j}}}\right)=:h(q_{\upsilon},n).$
It is straightforward that $h(q_{\upsilon},n)$ is increasing in both $q_{\upsilon}\geqslant 2$ and $n\geqslant 1$. In consequence, as $(q_{\upsilon},n)\neq(2,2)$, we either have $h(q_{\upsilon},n)\geqslant h(2,3)\approx 3.69$ or $h(q_{\upsilon},n)\geqslant h(3,2)\approx 17.75$. In any case, we arrived at the desired conclusion. ∎